A) Find the slope and y-intercept of 6x+10y=-20.
B)Write the equation of the line that passes through points (-2,18) with a slope of -5.
C)Write the equation of teh line with x-intercept of (16,0) and a y-intercept of (0-4)

Answer by stanbon(75887)

A) Find the slope and y-intercept of 6x+10y=-20.
y = (-3/5)x - 2
Slope = -3/5
intercept = -2
----------------------
B)Write the equation of the line that passes through points (-2,18) with a slope of -5.
y=mx+b where y=18 when x=-2 and m=-5
18 = -5*-2 + b
b = 8
EQUATION:
y = -5x +8
------------------------
C)Write the equation of the line with x-intercept of (16,0) and a y-intercept of (0-4)
Slope = [-4-0]/[0-16] = 1/4
Intercept = -4
EQUATION:
y = (1/4)x -4
=================
Cheers,
Stan H.